OpenStudy (anonymous):

What are the solutions & x-intercepts? 4x^2=12 Also Give exact and approximate solutions to three decimal places. y^2-6y+9=64

OpenStudy (anonymous):

First problem.\[4x^2=12 \implies x^2=3 \implies x= \pm \sqrt3\]

OpenStudy (anonymous):

Second problem. \[y^2-6y+9=64 \implies y^2-6y-55=(y-11)(y+5)=0 \implies y=11~or~ y=-5\]

OpenStudy (anonymous):

These are the exact answers; I leave the decimal approximation to you and your calculator.

OpenStudy (anonymous):

For the second problem, I solved it as if it were stated in the usual way, y=f(x). If we think that x=f(y), the x intercept is the constant term, i.e., x==55.